### v3

# Guan, Basciftci, and Van Hentenryck

[…]
Guan, Basciftci, and Van Hentenryck
[…]
Bilevel Optimization and Heuristics for Transit Network Design with Latent Demand
[…]
Bilevel Optimization and Heuristic Algorithms for Integrating Latent Demand
into the Design of Large-Scale Transit Systems
[…]
Capturing latent demand has a pivotal role in designing transit services as omitting these riders can lead to poor quality of service and/or additional costs. This paper explores this topic in the design of transit networks by considering the perspectives of both the transit agencies and riders. The paper presents a generic bilevel optimization model, namely the Transit Networks Design with Adoptions (TN-DA), that considers the network design decisions in the leader problem, and routing of the riders in the follower problem under the given network design, while allowing a black-box choice function for representing the adoption behavior of latent demand. The paper then identifies structural properties of the optimal solution of the TN-DA problem, which are desirable for transit agencies for capturing adoption behavior of the riders. The paper further provides guideline metrics for the transit agencies based on these desired adoption properties. Due to the computational complexity of this bilevel problem, the paper proposes five efficient heuristic algorithms to solve large-scale instances, which leverage an iterative procedure by solving a simpler version of the TN-DA problem and integrating the evaluation of rider choices. These algorithms either satisfy the desired properties of the optimal solution or provide fast approximations. The paper presents extensive large-scale case studies on two different transit systems by utilizing real datasets: (i) On-demand Multimodal Transit Systems (ODMTS) and (ii) Scooters-Connected Transit Systems (SCTS). The results demonstrate that the heuristic algorithms can find high-quality solutions much faster than exact approaches over various instances, while satisfying key adoption properties of the optimal solutions.
[…]
bilevel optimization, transportation network design, algorithm design, heuristics, latent demand, on-demand services
1 Introduction
Transit agencies have a critical concern in designing their transportation networks: how to ensure that a transit network design will capture ridership accurately and, in particular, customer adoption. This interaction between transit agencies (or in general, network designers) and latent riders, who represent the potential riders that can adopt the transit system, can be considered as a game in the context of the transit network design problem. Essentially, this problems involves decision-making by both transit agencies and latent riders with potentially conflicting objectives. The objective of the transit agency revolves around boosting ridership by attracting latent riders while containing the rise in operational expenses. Simultaneously, latent riders expect to access high-quality services, which might incur substantial cost increases for the transit agency.
The decisions made by the transit agencies and riders are usually studied separately. More specifically, transit agencies focus on designing networks with estimated travel demand, a subject extensively explored in literature (e.g., Borndörfer et al. (2007), Stiglic et al. (2018), Mahéo et al. (2019), Cats et al. (2025)). However, omitting potential riders that can adopt the transit system during the network design can lead to suboptimal designs with lower rider adoption and higher costs. As an example, this phenomena is observed in designing On-Demand Multimodal Transit Systems, which is an emerging form of public transportation system that tightly integrates on-demand shuttles with fixed transit services such as high-frequency buses and urban rail systems. Specifically, the transit agency tends to have higher costs and can provide poor quality of service for certain riders if the behavior of latent riders are omitted during the network design resulting in non-effective utilization of on-demand shuttles (Basciftci and Van Hentenryck 2020, 2023). When it comes to modeling the choices of the riders, research leans towards the utilization of logit and machine learning models for demand analysis (Xie et al. 2003, Hagenauer and Helbich 2017, Zhao et al. 2020). However, it remains critical to integrate rider choices to the transit network design problems while taking into account the decision making processes of both the transit agency and riders. Consequently, this paper explores the possibilities of integrating the aforementioned two pieces into one single framework through a game theoretic approach. Specifically, the paper proposes a bilevel optimization framework which is called Transit Network Design with Adoptions (TN-DA), with a rider choice modeling integrated as a black-box-like component. The optimal solution of the TN-DA problem can be viewed as the equilibrium point between the transit agency and the latent riders. Furthermore, the TN-DA framework is highly generalized and can be tailored to various network design problems.
Due to the bilevel nature of the problem with the combinatorial choice function of the latent riders, solving TN-DA can be practically challenging, especially for large-scale instances. This paper thus explores whether heuristics can provide fast approximations to the optimal solutions of large-scale instances.
To this end, this paper first identifies the structural properties of the optimal solution of the TN-DA problem, in terms of the adoption behavior. These properties are desirable by the transit agencies in establishing the equilibrium between the transit agency and riders. Then, the paper presents an iterative solution framework to significantly reduce the problem complexity by splitting the original bilevel framework into two components: (i) a regular transit network design problem without rider adoption and (ii) an evaluation component that uses a choice model to capture the adoption behavior and decides the set of riders and the network to be used in the next iteration. The heuristic algorithms combine these two components by increasing incrementally the sets of riders to be considered in the design and/or by progressively adding more fixed routes in the design, while guaranteeing to satisfy desired adoption properties. Therefore, this paper then proposes five heuristic algorithms to efficiently approximate the optimal solution of the TN-DA problem. The differences among these algorithms mostly occur at the transitions between these two primary steps, e.g, how they reshape the demand data and how they fix certain arcs in the network.
The effectiveness of these five heuristic algorithms is evaluated across two distinct transit network design problems. In addition to the objective value that can be used for evaluating the computational performance of the algorithms, this paper further proposes new metrics that assess the quality of the network designs by measuring the trade-off between the transit agency and the potential new users based on the properties of optimal solutions.
The primary contributions of this study can be summarized as follows:
  1. 1.
    This paper introduces the TN-DA framework, a generic bilevel optimization model that captures latent demand while designing transit network. The paper identifies adoption properties of optimal solutions, which inform heuristic design and new metrics: False Rejection Rate and False Adoption Rate, offering insights for evaluating and selecting transit network designs.
  2. 2.
    This paper proposes five heuristic algorithms to efficiently approximate optimal solutions of the TN-DA problem that can satisfy desired adoption properties or provide rapid approximations to the optimal solution.
  3. 3.
    This paper demonstrates the benefits of the heuristic algorithms on the TN-DA problem through large-scale case studies using real data on two transit systems: (i) On-demand Multimodal Transit Systems (ODMTS) and (ii) Scooter-Connected Transit Systems (SCTS). The results over various instances show that the heuristics are much more efficient than exact approaches by obtaining high-quality solutions in significantly shorter amounts of time while maintaining key adoption properties.
The remainder of this paper is organized as follows: Section 2 first presents the related work in transit network design and its connection with travel mode choices, followed by a discussion on the application of heuristic algorithms to various mobility problems. Section 3 introduces the generalized TN-DA problem. The five iterative algorithms are introduced in Section 4. Two detailed TN-DA examples, namely ODMTS-DA and SCTS-DA, are fully demonstrated with realistic case studies in Section 5 and 6. The conclusion and future research directions are discussed in Section 7.
2 Related Literature
Many existing methods for addressing transit network design problems are based on the assumption that the passenger demand is known and fixed (Schöbel 2012), whereas fewer number of studies have considered design/planning problems together with latent demand—potential customers who might be attracted by the proposed solutions. Klier and Haase (2008) presented an optimization model without a fixed demand by employing a linear demand function of the expected travel time to model its impact on transit demand. In another study, while maximizing the number of expected passengers, the same authors employed a logit model to estimate the mode choice between the best transit route and an alternative travel mode (Klier and Haase 2015). Additionally, a series of studies presented by Canca et al. comprehensively addressed this problem. Their first study introduced a model that simultaneously determines the network design, line planning, capacity, fleet investment, and passengers mode and route choices (Canca et al. 2016). Later on, two follow-up studies proposed adaptive large neighborhood search metaheuristics to solve variations of this model over large-scale data, respectively (Canca et al. 2017, 2019). Similarly, Laporte et al. (2005) aims to maximize trip coverage while considering rider behavior during the planning of rapid transit networks and Laporte et al. (2011) provide a review on this area. Although these studies highlight the importance of integrating rider behavior during the network design, it remains critical to consider the decision making of both the transit agency and riders within a multi-level formulation.
A bilevel formulation is a mathematical program where certain variables are constrained by the solution of another optimization problem, as described by Kleinert et al. (2021). This approach is valuable for modeling hierarchical decision processes and has extensive applications in transportation. Among the studies that consider latent ridership, Basciftci and Van Hentenryck (2020, 2023) introduced rider adoption awareness to the design of On-Demand Multimodal Transit System (ODMTS). These studies indicate that a bilevel optimization framework needs to be utilized to model the problem to ensure a fair design for the transit system that provides accessible transit services to the areas with existing demand and potential demand. The optimization model includes a personalized choice model which associates adoption choices with the time and cost of trips in the ODMTS. These two studies and a follow-up study (Guan et al. 2024) provide significant insights into the design of methods that solve the problem under different choice model assumptions. However, these methods still encounter computational challenges in efficiently solving large-scale problems. Building on these studies, this paper presents a generalized bilevel formulation applicable to a broad range of transit network design problems with latent demand by allowing more comprehensive choice and price models, while proposing heuristic algorithms providing desirable properties of the optimal designs with improved computational efficiency.
In general, iterative approaches are widely applied to the field of transportation science, especially for complex large-scale problems. For example, the following studies investigated iterative approaches on problems related to large-scale network design such as designing distribution networks and planning public transit for large urban areas (Gattermann et al. 2016, Schöbel 2017, Cipriani et al. 2012, Liu et al. 2022). Studies presented by Siebert and Goerigk (2013) and Polinder et al. (2022) demonstrate a typical method for employing iterative algorithms in network design, which involves designing a network, establishing timetables, and then iteratively refining the design. When expanding the network design problem from city-wise to country-wise, another example of using iterative approaches to solve large-scale public transit problems is the study presented by Liu et al. (2019). Similar to ODMTS that will be investigated in Section 5, the goal of the problem is to optimize both the operational costs and passenger travel convenience. However, instead of using time to directly measure travel convenience, a conversion from travel time to operational income is utilized. Two case studies using the Chinese high-speed rail data presented by the authors demonstrate that iterative algorithms can efficiently solve large-scale public transit planning problems while providing profitable solutions. Furthermore, iterative approaches are utilized to solve more general network design problems such as the incomplete hub location problems—selecting hub locations without the restriction of fully-connected networks (Dai et al. 2019) and the fleet assignment problem—assigning a fleet while maximizing expected revenue (Dumas et al. 2009). The latter study inspired Kroon et al. (2015), and they proposed a three-step iterative approach to optimize the rolling stock to the disrupted passenger flows. Although these studies do not explicitly address latent demand in network design and omit investigating properties of optimal solutions, they highlight the efficiency of using iterative algorithms in designing large-scale logistic and urban transit networks.
Besides designing public transit networks, iterative approaches are also applied to other topics in the field of transportation science when the original problem has a bilevel nature. Yu et al. (2010) first proposed a bilevel optimization model to determine the optimal bus frequencies to minimize the total passengers’ travel time, and an iterative algorithm that consists of a genetic algorithm and a label-marking method is then proposed to solve this model. Tawfik et al. (2022) provided another example of applying iterative algorithms on bilevel problems by continuing on a previous study where a bilevel model formulates service design and pricing for freight networks (Tawfik and Limbourg 2019). Lodi et al. (2016) studied game theoretic frameworks between the transit agency and bus operators for optimizing certain contractual relationships in public transit systems, and employed iterative schemes as part of their solution procedure.
So far, not much attention has been paid to the role of iterative algorithms when integrating public transit planning and passenger mode choices. One of the earliest approaches is presented by Lee and Vuchic (2005). They consider a transit network design problem with variable transit demand under a given fixed total demand that can be split into transit mode or personal vehicle mode, and this setting is also adopted by Fan and Machemehl (2006) within a genetic algorithm. The approaches used by these studies share two similarities with this paper—(i) the transit demand is variable such that riders can choose between transit services and personally-owned vehicles and (ii) at each iteration, the algorithm reshapes the transit demand then redesigns the transit network. On the other hand, this paper proposes a systematic approach for integrating rider preferences into the transit network design problem by introducing five iterative algorithms and leveraging the bilevel nature of the original problem. Three of these algorithms are driven by modifying trip sets considered in each iteration, whereas the other two are mainly driven by arc selection in the network. Furthermore, beside efficiently solving this large-scale problem, the iterative algorithms in this study also bring novelties by providing network designs with guaranteed adoption and rejection properties with insightful metrics that are practically useful to transit agencies.
3 Transit Network Design with Adoptions Problem (TN-DA)
This section introduces the general formulation of the Transit Network Design with Adoptions (TN-DA) problem. The motivation for the TN-DA problem stems from the desire to establish a transit network design $\mathbf{z}$ that concurrently serves the core riders (existing riders of the transit system) and captures latent demand with their adoption choice characteristics. Table 1 provides a summary of the nomenclatures that will be utilized in the remainder of this paper. Sections 3.1 and 3.2 describes the problem and introduces a generalized bilevel optimization framework for the TN-DA problem, respectively.
[…]
Symbol Definition
General Sets:
NN The set of all locations considered in the problem
$H$ The set of connecting points for the network, $H\subseteq N$
$T$ The set of all trips
$T_{core}$ The set of core trips, i.e., existing and fixed transit demand. Assumed to use transit under any conditions.
$T_{latent}$ The set of latent trips, i.e., latent demand. They are the potential customers of the system.
$\hat{T}$ The set of input trips for the TN-DFD problem. It is the union of $T_{core}$ and a subset of $T_{latent}$.
$M$ The set of possible modes, e.g., shuttles, buses, scooters, and walk.
$Z$ The set of design arcs that are considered in the design problem
$\mathcal{Z}$ The set of all feasible solutions
General Parameters:
$d_{ij}^{m}$ Travel distance between stops $i,j\in N$ under mode $m$
$t_{ij}^{m}$ Travel time between stops $i,j\in N$ under mode $m$
$p^{r}$ The number of riders in trip $r$
$g_{\pi}^{r}$ Travel cost for a rider in trip $r$ under path $\pi$
$\varphi_{\pi}^{r}$ Transit agency’s revenue from a rider in trip $r$ under path $\pi$
$\zeta_{\pi}^{r}$ Transit agency’s net-cost to serve a rider in trip $r$ under path $\pi$
$t_{\pi}^{r}$ Travel time for a rider in trip $r$ under path $\pi$
$t_{cur}^{r}$ Travel time of a latent trip $r\in T_{latent}$ with its current travel mode
$\alpha^{r}$ Adoption factor of a latent trip $r\in T_{latent}$, used in Equation (3)
Decision Variables:
$z_{hl}$ A binary variable indicating if a network arc between $h,l\in H$ is open
$\delta^{r}$ A binary variable indicating adoption decision of trip $r$, dynamically depending on $\mathbf{z}$
Sets for Algorithms:
$\overline{T}^{k}$ The set of trips considered by the algorithms at iteration $k$
$T_{adp}^{k}$ The set of latent trips that adopt network design $\mathbf{z}^{k}$ at iteration $k$
$C$ The set of all adopted trips in previous iterations (Algorithms $\rho$-GRAD and $\rho$-GAGR)
$T_{\rho}^{k}$ The set of latent trips that transferred to $C$ at iteration $k$ (Algorithms $\rho$-GRAD and $\rho$-GAGR)
$R$ The set of all rejected trips in previous iterations (Algorithms $\eta$-GRRE and $\rho$-GAGR)
$T_{m}^{k}$ The set of latent trips that transferred to $R$ at iteration $k$ (Algorithms $\eta$-GRRE and $\rho$-GAGR)
$D^{k}$ The set of candidate sub-networks at iteration $k$ (Algorithms arc-S1 and arc-S2)
$T_{next}^{k}$ The set of latent trips for next iteration’s expansion (Algorithms arc-S1 and arc-S2)
Parameters for Algorithms:
$\rho$ The amount of trips transferred to $C$ at each iteration (Algorithms $\rho$-GRAD and $\rho$-GAGR)
$\eta$ The step-size applied to $m$ (Algorithms $\eta$-GRRE and $\rho$-GAGR)
Variables for Algorithms:
$k$ An integer that records algorithms’ iterations
$m$ The amount of trips transferred to $R$ at each iteration (Algorithms $\eta$-GRRE and $\rho$-GAGR)
$B^{k}$ The upper bound of design objective at iteration $k$ (Algorithms arc-S1 and arc-S2)
$p$ Stage indicator (Algorithms arc-S1 and arc-S2)
[…]
Table 1: TN-DA Notations. (For simplicity and consistency with others, notations such as $t_{\pi^{r}}$ are written as $t_{\pi}^{r}$.)
3.1 Problem Description
The TN-DA problem revolves around a set of nodes, denoted by $N$, which represents all the locations involved in the problem. Among these nodes, a subset $H\subseteq N$ represents potential connecting points for network arcs, e.g., high-frequency buses in a transit system. A network is represented by $\mathbf{z}\in\{0,1\}^{|H|\times|H|}$, where the binary variable $z_{hl}$ represents whether an arc connecting nodes $h$ and $l$ in the set $H$ is open or not. For any nodes $i,j\in N$, $t_{ij}^{m}$ and $d_{ij}^{m}$ represent the time and distance under travel mode $m$, respectively, between these nodes, where $m\in M$. The set of modes $M$ can include various transportation modes considered in the problem, such as rail, bus, and walk.
To incorporate preferences of the riders into the planning process, the problem distinguishes between two sets of demand. The first set, denoted as $T_{core}$, represents riders who rely on transit systems and are guaranteed to use the newly designed system once it is deployed. A key assumption is that $T_{core}$ remains fixed, meaning all individuals within this set are always transit users. The second set, denoted as $T_{latent}$, consists of riders who currently use other types of transportation and have the option to switch to the new transit system. One can consider an example where a new transit system, denoted as $\mathbf{z}$, is proposed to replace the existing system. Core riders are individuals who currently utilize the existing transit system, and they are expected to continue using public transportation when the new system is implemented. On the other hand, latent riders do not currently rely on the existing transit systems and instead choose to drive their own cars for their travel needs. These individuals have not yet adopted public transportation as their preferred mode of travel. However, the introduction of a newly-designed transit network has the potential to attract and persuade latent riders to adopt public transit. The complete set of trips, denoted by $T$, is formed by combining $T_{core}$ and $T_{latent}$. Each trip, $r\in T$, consists of an origin location, $or^{r}\in N$, a destination location, $de^{r}\in N$, and a group of $p^{r}$ riders who travel together, where $p^{r}\in\mathbb{Z}{+}$. Within the context of this paper, the term “trip rr” denotes a set of $p^{r}$ riders who undertake a journey collectively from their origin to destination. This implies that they share the same routes and possess common attributes and characteristics.
A path $\pi^{r}$ in the transit network for a given trip $r$ can be dynamically constructed based on the state of the network. The path should start at the trip origin node ${or}^{r}$, end at the trip destination node ${de}^{r}$, and may include zero or multiple intermediate nodes in between. The trip duration, cost, and revenue for a single rider to complete path $\pi^{r}$ can be calculated by considering the legs of the path. The duration, cost, and revenue values are represented as $t_{\pi}^{r}$, $g_{\pi}^{r}$, $\varphi_{\pi}^{r}$, respectively, and the net-cost $\zeta_{\pi}^{r}$ incurred by the transit agency for serving the path $\pi^{r}$ can be subsequently computed as $\zeta_{\pi}^{r}=g_{\pi}^{r}-\varphi_{\pi}^{r}$. It is important to note that in the given context, the terms cost, revenue, and net-cost do not necessarily refer to the monetary value associated with serving trip $r$. Instead, they represent customizable measures that can be defined based on specific criteria or considerations, although the most common approach is to use monetary measures.
3.2 Bilevel Optimization Framework
Figure 1 presents a generalized bilevel optimization framework for capturing the travel mode adoption during the transit network design by considering the core riders ($T_{core}$) and potential riders ($T_{latent}$) that can adopt the new system. The leader problem (1) corresponds to the network operator, which aims at designing a transit network, i.e., determining which arcs are connected to each other under adoption and problem specific considerations. The objective of the leader problem (1a) minimizes the sum of (i) the investment cost of opening network arcs, which is denoted by a function $\text{\sc Invest}(\mathbf{z})$, (ii) the net-cost incurred by core riders, and (iii) the net-cost of latent riders that are adopting the new transit system. Constraint (1b) corresponds to the customized restrictions in constructing the transit network design. Additional constraints on the arc variables $\mathbf{z}$ can be introduced based on specific problems; in this section, only the generic form is presented. Constraint (1c) represents the mode choice function of a latent trip depending on path $\pi^{r}$.
$\min_{\mathbf{z},\delta^{r}}\quad$ $\text{\sc Invest}(\mathbf{z})+\sum_{r\in T_{core}}p^{r}(g^{r}-\varphi^{r})+\sum_{r\in T_{latent}}p^{r}\delta^{r}(g^{r}-\varphi^{r})$ (1a)
s.t. $\mathbf{z}\in\mathcal{Z}\subseteq\{0,1\}^{|H|\times|H|}$ (1b)
$\delta^{r}=\mathcal{C}^{r}(\pi^{r})\quad\forall r\in T_{latent}$ (1c)
$\delta^{r}\in\{0,1\}\quad\forall r\in T_{latent}$ (1d)
where $(\pi^{r},g^{r},\varphi^{r})$ are a solution to the optimization problem $\text{\sc OPT-PATH}^{r}(\mathbf{z})$ as follows:
[…]
$g^{r}=$ $\text{\sc PATH-COST}(\pi^{r})\quad\forall r\in T$ (2b)
$\varphi^{r}=$ $\text{\sc PATH-REVENUE}(\pi^{r})\quad\forall r\in T$ (2c)
Figure 1: The Generalized Optimization Model for the Transit Network Design with Adoption (TN-DA) Problem.
Adoption Choice Model
In the TN-DA problem, a choice model is employed to capture the adoption decision of latent riders, taking a path $\pi$ as its input. The binary output of $\mathcal{C}^{r}$ (i.e., $\delta^{r}$ in (1a) and (1c)) indicates whether all riders in latent trip $r$ will adopt or reject the transit network design $\mathbf{z}$. The choice model $\mathcal{C}^{r}$ can be arbitrarily complex, which can range from a simple decision model to a sophisticated machine learning-based model. If available, information such as geographical and demographic data of riders can be incorporated into $\mathcal{C}^{r}$. The crucial aspect is that $\mathcal{C}^{r}$ is treated as a black-box, which implies the generalizability of the proposed framework.
A simple example of such a model is presented in Equation (3), namely Time-Based. This particular model compares the trip duration $t_{\pi}^{r}$ with the travel duration $t_{cur}^{r}$ under the current travel mode of the riders, multiplied by a parameter $\alpha^{r}\geq 1$. Thus, if the travel duration under $\pi^{r}$ is less than the desired threshold of the riders, then this choice function returns 1, representing adoption of the designed system.
$\text{{\sc Time-Based}}:{\cal C}^{r}(\pi)\equiv\mathbbm{1}(t_{\pi}^{r}\leq\alpha^{r}\ t^{r}_{cur})$ (3)
Subproblem
[…]
To this end, the intuition behind the bilevel model can be summarized as follows: The network designer first obtains the design by considering the net-cost of the trips considered. Secondly, the follower problems determine the paths for the riders during operations. The choice function in the leader problem determines, for each trip with a mode choice, whether they would adopt the proposed path or elect to use their previous travel mode. If such riders decide to use the new system, the operator receives a revenue for that trip.
An instance of the TN-DA problem is the On-demand Multimodal Transit System Design with Adoptions (ODMTS-DA) problem, which was examined in previous studies by Basciftci and Van Hentenryck (2020, 2023), Guan et al. (2024). After introducing the heuristic algorithms in the subsequent Section 4, Section 5 will revisit the ODMTS-DA problem and explore its connection to the algorithms in details. Another example of the TN-DA problem, namely the Scooters-Connected Transit Systems Design with Adoptions (SCTS-DA) problem, will be showcased in Section 6. Both case studies concentrate on emerging technologies in the evolving landscape of public transportation.
4 Heuristic Algorithms and Design Properties
Section 3 presents the generalized TN-DA problem within a bilevel optimization framework. However, the presence of hierarchical decision processes in bilevel optimization problems, combined with the combinatorial nature of the TN-DA problem with rider choices, introduce additional complexities, making it inherently difficult to solve in general. Therefore, this section presents heuristic algorithms to find high-quality solutions of the TN-DA problem in reasonable time, especially for large-scale instances. The heuristic algorithms use, as a key component, the Transit Network Design with Fixed Demand (TN-DFD) problem which is presented in Figure 2. Compared to the TN-DA, the TN-DFD removes the choice function constraint (1c) and the decision variables of the form $\delta^{r}$ that correspond to the user choices. The heuristic algorithms will use TN-DFD with a set of trips $\hat{T}\subseteq T$ that is always constructed from all core trips and a subset of the latent trips.
$\min_{\mathbf{z}}\quad$ $\text{\sc INVEST}(\mathbf{z})+\sum_{r\in\hat{T}}p^{r}(g^{r}-\varphi^{r})$ (4a)
s.t. (1b)
where $(\pi^{r},g^{r},\varphi^{r})$ are a solution to the optimization problem $\text{\sc OPT-PATH}^{r}(\mathbf{z})$ as follows:
[…]
Figure 2: The Generalized Formulation for the Transit Network Design with Fixed Demand $\hat{T}$: TN-DFD($\hat{T}$).
By leveraging the TN-DFD problem, the overall idea of the heuristics can be summarized in Algorithm 1. All algorithms start their first iteration by solving the TN-DFD problem with the core trip set $T_{core}$ to obtain an initial transit network design, unless specified otherwise. Then, the set $\overline{T}^{k}$ denotes the set of trips considered by the algorithm at iteration $k$, i.e., $\hat{T}=\overline{T}^{k}$ at iteration $k$, where the latent trips are gradually added to $\overline{T}^{k}$ over the iterations depending on their adoption behavior under the resulting designs, while the trips in the core trip set remain in $\overline{T}^{k}$ as $k$ increases. In certain algorithms, network design can be constructed incrementally by identifying a specific set of arcs to be fixed at each iteration.
[…]
Algorithm 1 TN-DA Heuristic (Generic Form)
1: Set $k=0$, $\overline{T}^{0}=T_{core}$, no arcs are fixed, unless other input values are mentioned
2:while Stop Criteria Not Met do
3:  Solve $\mathbf{z}^{k}=\text{\sc TN-DFD}(\overline{T}^{k})$.
4:  Evaluate the design $\mathbf{z}^{k}$ and each latent trip $r\in T_{latent}$’s adoption decision $\delta^{r}$ using $\mathcal{C}^{r}$.
5:  Construct $\overline{T}^{k+1}$ using $T_{latent}$ and/or fix a specific set of arcs by applying certain rules while ensuring that $T_{core}\subseteq\overline{T}^{k+1}$.
6:  Update $k$.
7:end while
8:return $\mathbf{z}^{k}$ unless another design is more desirable under certain rules.
[…]
Property 1 (Correct Rejection)
[…]
Property 2 (Correct Adoption)
[…]
When a design satisfies these two properties, it represents an equilibrium between the transit agency and the latent riders. The following proposition links the optimal solution $\mathbf{z}^{*}$ of TN-DA and Properties 1 and 2.
Proposition 1
Assume $\mathbf{z}^{*}$ is the optimal solution of the TN-DA problem (Figure 1). There exists an input trip set $\hat{T}^{*}\subseteq T$ such that $\mathbf{z}^{*}=\mbox{{\sc TN-DFD}}(\hat{T}^{*})$ and $\mathbf{z}^{*}$ satisfies both Correct Rejection and Correct Adoption properties (Properties 1 and 2).
Proof.
[…]
The goal of the heuristic algorithms is to find the set $\hat{T}^{*}$, which is computationally challenging to obtain. Thus, the proposed algorithms aim at achieving one of two different goals: (1) efficiently finding an approximation of $\hat{T}^{*}$ that satisfies the Correct Rejection property or Correct Adoption property, or (2) rapidly finding an approximation to $\mathbf{z}^{*}$ regardless of these properties. The rest of this section is organized as follows. Section 4.1 introduces key concepts for evaluating the designs found by the heuristic algorithms. Sections 4.2 and 4.3 describe the trip-based iterative algorithms and arc-based iterative algorithms, respectively. The discussion of the two properties is also presented throughout Sections 4.2 and 4.3 as the iterative algorithms are introduced.
4.1 Key Concepts for Design Evaluation
This section summarizes multiple common features shared by all proposed iterative algorithms. Since the TN-DFD is employed to approximate the optimal solution of the TN-DA, the objective function (1a) of the TN-DA is utilized to evaluate any network design $\mathbf{z}$ produced by the TN-DFD, using Equation (6). Regardless of the set $\hat{T}$ used to obtain the design $\mathbf{z}$, the input trip set to evaluate $\mathbf{z}$ is always the full trip set $T$. Under the design $\mathbf{z}$, this evaluation is performed using the choice decisions $\delta^{r}$ from $\mathcal{C}^{r}$ and the optimal path $\pi^{r}$ associated with each trip $r$.
$\text{\sc Eval}(\mathbf{z})=\text{\sc Invest}(\mathbf{z})+\sum_{r\in T_{core}}p^{r}(g_{\pi}^{r}-\varphi_{\pi}^{r})+\sum_{r\in T_{latent}}p^{r}\delta^{r}(g_{\pi}^{r}-\varphi_{\pi}^{r})$ (6)
To evaluate the quality of the network designs found by the iterative algorithms, two additional metrics, i.e., False Rejection Rate ($\%R_{false}$) and False Adoption Rate ($\%A_{false}$), are introduced to measure the violation of the Correct Rejection and Correct Adoption properties (Properties 1 and 2). These two metrics are defined as follows:
$\%R_{false}=\cfrac{|\{r\in T_{latent}\setminus\hat{T}:r\text{ adopts design }\mathbf{z}\}|}{|T_{latent}|}\times\%100$
$\%A_{false}=\cfrac{|\{r\in T_{latent}\cap\hat{T}:r\text{ rejects design }\mathbf{z}\}|}{|T_{latent}|}\times\%100$
Here, $\%R_{false}$ indicates the percentage of latent trips that are excluded during the network design although they adopt the resulting design, implying false rejection, whereas $\%A_{false}$ corresponds to the percentage of latent trips that are included during the network design although they reject the resulting design, implying false adoption. Core trips are excluded from the calculation of these metrics, because they are always included in $\hat{T}$ and not associated with a choice function. The subsequent two Remarks highlight that obtaining zero False Rejection Rate or False Adoption Rate is neither a necessary nor a sufficient condition for finding an optimal solution of TN-DA as these conditions depend on the trip set considered in the TN-DFD problem. Nevertheless, they establish desired adoption properties by identifying an equilibrium between the transit agency and latent riders based on the latent riders considered during the network design.
Remark 1
It is possible that the solution of the TN-DFD with a trip set $\hat{T}\neq\hat{T}^{*}$ is a design $\mathbf{z}\neq\mathbf{z^{*}}$ that satisfies both Correct Rejection property and Correct Adoption property. Thus, satisfying both properties under an input trip set $\hat{T}$ does not necessarily imply that $\mathbf{z}$ is the optimal solution of TN-DA.
Remark 2
It is possible that the optimal solution of the TN-DFD with a trip set $\hat{T}\neq\hat{T}^{*}$ is the optimal design $\mathbf{z}^{*}$ of the TN-DA. In other words, $\mathbf{z}^{*}$ may be discovered by a trip set $\hat{T}\neq\hat{T}^{*}$ with $\%R_{false}\neq 0$ or $\%A_{false}\neq 0$ under $\hat{T}$.
4.2 Trip-based Iterative Algorithms
This section introduces the first category of iterative algorithms—the trip-based algorithms. In general, each iteration of this type of algorithm can be summarized by three fundamental steps. The first step solves the TN-DFD problem with $\hat{T}=\overline{T}^{k}$. The second step evaluates the design with $\text{\sc Eval}(\mathbf{z})$ (Equation (6)). The third step computes the new $\overline{T}^{k+1}$ for the next iteration. All trip-based iterative algorithms have identical first and second steps, with the only difference occurring in the third step. At the third step, the trip-based algorithms need to compute a customizable value for each trip $r$, denoted as $v^{r}$ (such as the net-cost $\zeta_{\pi}^{r}=g_{\pi}^{r}-\varphi_{\pi}^{r}$), based on the optimal path of that trip under the given network design. This value is then utilized in constructing the trip set to be considered in the subsequent iterations.
Three trip-based iterative algorithms are proposed. The first two are presented in Sections 4.2.1 and 4.2.2. Another one combining the previous two algorithms is explained in Section 4.2.3, which enables to explore more intermediate network designs, despite of its much longer expected running time. Section 4.2.4 presents some properties of the proposed trip-based algorithms.
4.2.1 Greedy Adoption Algorithm ($\rho$-GRAD)
The first trip-based iterative algorithm starts with the core trips and greedily adds adopting trips at each iteration. Algorithm $\rho$-GRAD maintains a set $C$ of trips and transfers latent trips from $T_{latent}$ to $C$ at each iteration. Specifically, at a given iteration $k$, the solution $\mathbf{z}^{k}$ identifies a set of adopting trips $T_{adp}^{k}\in T_{latent}$. It transfers $\rho$ trips from $T_{adp}^{k}$ to $C$, choosing those with the least $v^{r}$ values. This set of $\rho$ trips is denoted by $T_{\rho}^{k}$. The algorithm terminates when there are no latent trips to be absorbed by set $C$ at the end of an iteration, i.e., $|T_{\rho}^{k}|=0$. Algorithm $\rho$-GRAD satisfies the following property.
Proposition 2
The last design $\mathbf{z}^{k}$ found by Algorithm $\rho$-GRAD with trip set $\hat{T}=\overline{T}^{k}$ satisfies the Correct Rejection property (Property 1), i.e., $\%R_{false}=0$.
Proof.
[…]
Algorithm 2 $\rho$-GRAD
1: Set $k=0$, $\overline{T}^{0}=T_{core}$, $C=\emptyset$, unless other input values are specified.
2:Require: step-size parameter $\rho$
3:while True do
4:  Set $\mathbf{z}^{k}=\mbox{{\sc TN-DFD}}(\overline{T}^{k})$.
5:  Set $T_{check}=T_{latent}\setminus C$.
6:  for all $r\in T_{check}$ do
7:   Set $\pi^{r}$ = $argmin\text{\sc OPT-PATH}^{r}(\mathbf{z}^{k})$.
8:   Set $\delta^{r}={\cal C}^{r}(\pi^{r})$.
9:   if $\delta^{r}=1$ then
10:    Compute $v^{r}$ for trip $r$.
11:   end if
12:  end for
13:  Set $T_{adp}^{k}=\{r\in T_{check}\mid\delta^{r}=1\}$.
14:  Set $T_{\rho}^{k}$ as the set of $\rho$ trips from $T_{adp}^{k}$ with the lowest $v^{r}$.
15:  if $T_{\rho}^{k}=\emptyset$ then
16:   break.
17:  else
18:   Set $C=C\cup T_{\rho}^{k}$.
19:   Set $\overline{T}^{k+1}=T_{core}\cup C$.
20:   Set $k=k+1$.
21:  end if
22:end while
23:return $\mathbf{z}^{k}$
4.2.2 Greedy Rejection Algorithm ($\eta$-GRRE)
Algorithm $\eta$-GRRE greedily excludes trips from $T_{latent}$ and introduces a growing set $R$ of rejecting trips. In addition, a variable $m$ is used to indicate how many adopting trips from $T_{latent}\setminus R$ are included in the set $\overline{T}^{k}$ and $m$ is increased by $\eta$ at each iteration. Using $T_{latent}\setminus R$ directly would lead to many rejections, so the algorithm proceeds carefully when choosing $\overline{T}^{k}$ early on. Algorithm $\eta$-GRRE terminates when it meets the following two conditions at an iteration $k\geq 2$:
  1. 1.
    the design is stable, i.e., $\mathbf{z}^{k}=\mathbf{z}^{k-1}$;
  2. 2.
    $m-\eta\geq|T_{adp}^{k}|$.
Note that condition 1 alone is not adequate. Indeed, in early stages, the algorithm may tend to find the same design because $m$ is still relatively small. As a result, condition 2 is firstly employed to guarantee that the algorithms delve into a greater number of iterations. Nevertheless, it is crucial to acknowledge that beyond the point where $m$ exceeds $|T_{adp}^{k}|+\eta$, the algorithm will commence producing designs similar to those from earlier iterations. This is likely due to the decrease in $\hat{T}$ over iterations when many trips are included in $R$, which might also lead to oscillating results between iterations. Thus, condition 2 serves the purpose of striking a balance between these two effects. The designs found by Algorithm $\eta$-GRRE may not satisfy either the Correct Rejection property or the Correct Adoption property. Instead, this algorithm is particularly designed to provide a rapid approximation.
Remark 3
Because of possible oscillations in network designs between iterations, the $\eta$-GRRE algorithm reports the design with the minimal objective instead of the last design. These oscillations occur as the set $\overline{T}^{k}$ does not monotonically increase. Note that if the last design $\mathbf{z}^{k}$ is reported, then the algorithm satisfies the Correct Adoption property (Property 2) given that $\mathbf{z}^{k}=\mathbf{z}^{k-1}$.
[…]
Algorithm 3 $\eta$-GRRE
1: Set $k=0$, $m=0$, $R=\emptyset$, $\overline{T}^{0}=T_{core}$, and $o^{*}=\infty$, unless other input values are specified.
2:Require: step-size parameter $\eta$
3:while True do
4:  $\mathbf{z}^{k}=\mbox{{\sc TN-DFD}}(\overline{T}^{k})$.
5:  if $\text{\sc Eval}(\mathbf{z}^{k})<o^{*}$ then
6:   Set $o^{*}=\text{\sc Eval}(\mathbf{z}^{k})$.
7:   Set $\mathbf{z}_{min}=\mathbf{z}^{k}$.
8:  end if
9:  Set $T_{check}=T_{latent}\setminus R$.
10:  for all $r\in T_{check}$ do
11:   Set $\pi^{r}$ = $argmin\text{\sc OPT-PATH}^{r}(\mathbf{z}^{k})$
12:   Set $\delta^{r}={\cal C}^{r}(\pi^{r})$.
13:   if $\delta^{r}=1$ then
14:    Compute $v^{r}$ for trip $r$.
15:   else
16:    Set $R=R\cup\{r\}$.
17:   end if
18:  end for
19:  Set $T_{adp}^{k}=\{r\in T_{check}\mid\delta^{r}=1\}$.
20:  Set $m=m+\eta$.
21:  Set $T_{m}^{k}=$ the set of $m$ trips from $T_{adp}^{k}$with the lowest $v^{r}$.
22:  if the termination criteria is met then
23:   break.
24:  else
25:   Set $\overline{T}^{k+1}=T_{core}\cup T_{m}^{k}$;
26:   Set $k=k+1$.
27:  end if
28:end while
29:return $\mathbf{z}_{min}$
4.2.3 Greedy Adoption Algorithm with Greedy Rejection Subproblem ($\rho$-GAGR)
Algorithm $\rho$-GAGR is the same as Algorithm $\rho$-GRAD, except that it uses Algorithm $\eta$-GRRE as its subroutine instead of solving the TN-DFD problem to obtain a network design. This design decision makes it possible for Algorithm $\rho$-GAGR to explore many more designs than Algorithm $\rho$-GRAD. Algorithm $\rho$-GAGR shares the same terminating criteria with Algorithm $\rho$-GRAD. The full algorithm is presented in Appendix 8.
4.2.4 Additional Properties of Trip-based Algorithms
This section analyzes the change in the net-cost of the trips, as different sets of trips are considered during the network design. To this end, recall that $\zeta^{r}(\mathbf{z})=g^{r}(\mathbf{z})-\varphi^{r}(\mathbf{z})$ is defined as the optimum net-cost of the trip $r$ under the network design $\mathbf{z}$.
Proposition 3
Let $\mathbf{z}^{1}=\mbox{{\sc TN-DFD}}(T^{1})$ and $\mathbf{z}^{2}=\mbox{{\sc TN-DFD}}(T^{2})$ with $T^{1}\subseteq T^{2}$. Then $\sum\limits_{r\in T^{2}\setminus T^{1}}p^{r}(\zeta^{r}(\mathbf{z}^{2})-\zeta^{r}(\mathbf{z}^{1}))\leq 0$.
Proof.
Proof: Note that the optimum objective function value of the network design $\mathbf{z}^{2}$ under the set of trips $T^{2}$ can be written as $\sum\limits_{h,l\in H}\beta_{hl}\mathbf{z}^{2}_{hl}+\sum\limits_{r\in T^{2}}p^{r}\zeta^{r}(\mathbf{z}^{2})$. Since $\mathbf{z}^{1}$ is a feasible network design for this problem, the following relationship holds:
$\sum_{h,l\in H}\beta_{hl}\mathbf{z}^{2}_{hl}+\sum_{r\in T^{2}}p^{r}\zeta^{r}(\mathbf{z}^{2})\leq\sum_{h,l\in H}\beta_{hl}\mathbf{z}^{1}_{hl}+$ $\sum_{r\in T^{2}}p^{r}\zeta^{r}(\mathbf{z}^{1})$
$\sum_{h,l\in H}\beta_{hl}\mathbf{z}^{2}_{hl}+\sum_{r\in T^{1}}p^{r}\zeta^{r}(\mathbf{z}^{2})+\sum_{r\in T^{2}\setminus T^{1}}p^{r}\zeta^{r}(\mathbf{z}^{2})\leq\sum_{h,l\in H}\beta_{hl}\mathbf{z}^{1}_{hl}+$ $\sum_{r\in T^{1}}p^{r}\zeta^{r}(\mathbf{z}^{1})+\sum_{r\in T^{2}\setminus T^{1}}p^{r}\zeta^{r}(\mathbf{z}^{1})$
$\sum_{h,l\in H}\beta_{hl}(\mathbf{z}^{2}_{hl}-\mathbf{z}^{1}_{hl})+\sum_{r\in T^{1}}p^{r}(\zeta^{r}(\mathbf{z}^{2})-\zeta^{r}(\mathbf{z}^{1}))+$ $\sum_{r\in T^{2}\setminus T^{1}}p^{r}(\zeta^{r}(\mathbf{z}^{2})-\zeta^{r}(\mathbf{z}^{1}))\leq 0$
Following an analogous relationship by comparing the network designs under the set of trips $T^{1}$ results in
$\sum_{h,l\in H}\beta_{hl}\mathbf{z}^{2}_{hl}+\sum_{r\in T^{1}}p^{r}\zeta^{r}(\mathbf{z}^{2})$ $\geq\sum_{h,l\in H}\beta_{hl}\mathbf{z}^{1}_{hl}+\sum_{r\in T^{1}}p^{r}\zeta^{r}(\mathbf{z}^{1})$
$\sum_{h,l\in H}\beta_{hl}(\mathbf{z}^{2}_{hl}-\mathbf{z}^{1}_{hl})+\sum_{r\in T^{1}}p^{r}(\zeta^{r}(\mathbf{z}^{2})-\zeta^{r}(\mathbf{z}^{1}))$ $\geq 0$
[…]
Corollary 1
If the set $T^{2}\setminus T^{1}$ is a singleton $\{r\}$, then $\zeta^{r}(\mathbf{z}^{2})\leq\zeta^{r}(\mathbf{z}^{1})$.
These results have interesting consequences for Algorithms $\rho$-GRAD and $\rho$-GAGR. With a step size of 1, the added trip will not have a net-cost in the last design worse than those in the prior iteration. With larger step size, at least one of the added trips will satisfy this property. However, this property is not guaranteed for all the added trips in that iteration. Note that, as opposed to Algorithms $\rho$-GRAD and $\rho$-GAGR, this property does not necessarily hold for Algorithm $\eta$-GRRE as $\overline{T}^{k}\subseteq\overline{T}^{k+1}$ is not ensured in any iteration $k$.
4.3 Arc-based Iterative Algorithms
This section introduces the second type of iterative algorithms—the arc-based iterative algorithms. These algorithms begin with a pre-defined transit network design, denoted as $\mathbf{z}_{fixed}$, which can be either an empty network or an existing network such as a well-developed rail system. At the beginning of each iteration, these algorithms solve the TN-DFD problem and obtain a network design building on the open arcs in $\mathbf{z}_{fixed}$. Then, a sub-network from the resulting design is incrementally added in each iteration to enhance the overall network. The input trip set considered for the TN-DFD problem is then expanded, considering the adoption decisions of the riders. These steps are iterated until no improving sub-network can be added to the design. Since these algorithms select the best sub-network with respect to the improvement in the overall objective, arc-based algorithms produce a sequence of network designs whose evaluated objectives are monotonically decreasing. As a result, they find primal solutions quickly, which is not guaranteed in trip-based algorithms. In summary, arc-based algorithms satisfy the following properties:
  1. 1.
    The design objective, computed by function $\text{\sc Eval}(\mathbf{z})$ (Equation (6)), monotonically decreases.
  2. 2.
    The transit network design $\mathbf{z}_{fixed}$ monotonically increases.
This paper proposes two arc-based algorithms. Section 4.3.1 presents the Arc-based Greedy Algorithm arc-S1, and Section 4.3.2 presents the algorithm arc-S2 that extends arc-S1 to overcome the potential challenges faced by arc-S1 in terms of run time and solution quality. Both algorithms are parametrized by one or more functions that expand the set of users at each iteration. Section 4.3.3 introduces expansion rules to construct these Expand functions used by the algorithms.
4.3.1 Arc-based Greedy Algorithm (arc-S1)
[…]
Algorithm 4 arc-S1
1: Set $k=0$, $\mathbf{z}_{fixed}=\vec{0}$, $\overline{T}^{0}=T_{core}$, $B^{0}=\infty$.
2:Require: Expand function for increasing the set of users considered.
3:while True do
4:  Set $\mathbf{z}_{temp}^{k}$ = TN-DFD($\overline{T}^{k}$$\{(h,l)\mid{{z}_{fixed}}_{hl}=1\ \>\ h,l\in H\}$).
5:  Set $\mathbf{z}^{k}_{unfixed}=\mathbf{z}_{temp}^{k}-\mathbf{z}_{fixed}$.
6:  Set $D^{k}$ to the set of all candidate sub-networks in $\mathbf{z}^{k}_{unfixed}$.
7:  if $D^{k}=\emptyset$ then
8:   break.
9:  end if
10:  for all $\mathbf{d}\in D^{k}$ do
11:   if $(\mathbf{z}_{fixed}+\mathbf{d})\in\mathcal{Z}$ (i.e. if the design is feasible) then
12:    Set $obj_{\mathbf{d}}=\text{\sc Eval}(\mathbf{z}_{fixed}+\mathbf{d})$.
13:   end if
14:  end for
[…]
16:  if ${obj}^{k}<B^{k}$ then
17:   Set $B^{k+1}={obj}^{k}$.
18:  else
19:   break.
20:  end if
21:  Set $\mathbf{z}_{fixed}$ = $\mathbf{z}_{fixed}+\mathbf{d}^{k}$.
22:  Set $\mathbf{z}^{k}$ = $\mathbf{z}_{fixed}$
23:  Set $T_{next}^{k}$ = Expand($T_{latent},\mathbf{z}^{k}$).
24:  Set $\overline{T}^{k+1}=\overline{T}^{k}\cup T_{next}^{k}$.
25:  Set $k=k+1$.
26:end while
27:return $\mathbf{z}^{k}$
Algorithm arc-S1 describes the arc-based greedy algorithm. It maintains an upper bound $B^{k}$ that represents the best design objective found up to iteration $k$. Algorithm arc-S1 also uses a generalization of the TN-DFD, where the second parameter represents the set of arcs that must be fixed to 1 in the TN-DFD solution (line 4).
During each iteration, after solving such a generalized TN-DFD, the resulting design $\mathbf{z}_{temp}^{k}$ is decomposed into two parts: $\mathbf{z}_{fixed}$ and $\mathbf{z}_{unfixed}^{k}$ (line 5). The partial design $\mathbf{z}_{fixed}$ denotes all the arcs which were constrained to be opened in the TN-DFD, while $\mathbf{z}^{k}_{unfixed}$ is the remaining assignment in $\mathbf{z}_{temp}^{k}$. Then, the sub-network candidate pool $D^{k}$ needs to be carefully determined while ensuring that the addition of sub-networks does not lead to infeasibility in future iterations by considering the specific characteristics and requirements of the problem at hand. In particular, each iteration attempts to add a new sub-network $\mathbf{d}^{k}\in D^{k}$ to the fixed design $\mathbf{z}_{fixed}$. Algorithm arc-S1 does this greedily, adding the sub-network $\mathbf{d}^{k}$ which provides the greatest decrease in $B^{k}$ (lines 10–18). Therefore, the series $B^{0},\ldots,B^{k}$ decreases monotonically over the iterations, while the design $\mathbf{z}_{fixed}$ expands (line 21), where the computational time of each iteration tends to decrease as the number of variables decreases. Algorithm arc-S1 also expands the trips to consider in the TN-DFD (lines 23–24). The algorithm terminates when the design objective bound is not decreasing, i.e., $B^{k}=B^{k+1}$, which can happen under the following two cases:
  1. 1.
    $D^{k}$ is empty, i.e., it does not contain any new candidate sub-networks.
  2. 2.
    No sub-networks in $D^{k}$ provides a better objective value.
4.3.2 Arc-based Extended Greedy Algorithm (arc-S2)
The Expand function used in Algorithm 4 can have a significant impact on the performance of the algorithm, as it provides the balance in avoiding the set of trips in $\overline{T}^{k}$ from growing too quickly or too slowly over the iterations. A rapid expansion of the set $\overline{T}^{k}$ can introduce computational challenges when solving the TN-DFD problem. Conversely, a slow expansion of $\overline{T}^{k}$ may lead to premature termination of the algorithm, resulting in poor solutions. To overcome this difficulty, arc-S2 implements a two-phase extension of arc-S1 by using two Expand functions. Initially, arc-S2 expands the set of trips slowly until convergence using function Expand1. Then arc-S2 enters its second phase where the trip set increases faster. Both phases implement arc-S1, except that the second phase starts with the fixed design and the set of trips produced by phase 1. The full algorithm is presented in Appendix 8.
4.3.3 Trip Expansion Rules
The arc-based algorithms have been examined with various expansion rules to build the Expand functions over different transit systems. The initial rule, referred to as the “Full Adoption Expansion Rule”, considers all latent trips that adopt $\mathbf{z}^{k}$ to be included in the subsequent iteration. This rule is intuitive and can be generally applied to various types of transit systems. In the following Sections 5 and 6, the paper will further demonstrate how expansion rules can be carefully designed within a specific context by leveraging the underlying problem. This will showcase the flexibility of the approach in adapting to different transit system scenarios while maintaining the underlying principles discussed earlier. When using this expansion rule, Algorithm arc-S1 satisfies the Correct Rejection property. This result extends to Algorithm arc-S2.
Proposition 4
The transit network design returned by Algorithm arc-S1 with Full Adoption Expansion Rule satisfies the Correct Rejection property (Property 1), i.e., $\%R_{false}=0$.
Proof.
[…]
Corollary 2
The transit network design returned by Algorithm arc-S2 with the Full Adoption Expansion Rule in the second phase satisfies the Correct Rejection property (Property 1).
5 TN-DA Example 1: ODMTS Design with Adoptions Problem
This section presents a detailed example of the TN-DA problem with three realistic case studies, focusing on a specific type of transit network known as On-demand Multimodal Transit Systems (ODMTS). Therefore, within this particular setting, the TN-DA problem is referred to as ODMTS-DA (ODMTS Design with Adoptions) Problem.
As depicted in Figure 3, ODMTS are operated around a number of hubs that serve as stops for buses and rail. The on-demand shuttles are primarily employed as the feeders to the hubs and represent an effective solution to the first-and-last-mile problem faced by most of the transit agencies. Several simulation studies have demonstrated that ODMTS can provide considerable cost and convenience benefits in different cities under diverse circumstances (Mahéo et al. 2019, Dalmeijer and Van Hentenryck 2020, Auad et al. 2021). In practice, ODMTS were also successfully deployed in the cities of Atlanta (Van Hentenryck et al. 2023) and Savannah (Chatham Area Transit 2024) in 2022 and 2024—2025, respectively. Using the ODMTS-DA as an example of the TN-DA, the goal of this section is to demonstrate the practical advantages and outcomes of the heuristic algorithms in a real-world transit network design problem.
Figure 3: Picture that illustrates the On-demand Multimodal Transit Systems (ODMTS). The solid circles represent transit hubs. The solid line represents a typical ODMTS path offered by the transit agency.
In the context of ODMTS-DA, two previous studies addressed this problem by modeling it with bilevel optimization frameworks (Basciftci and Van Hentenryck 2020, 2023). In particular, the agency proposes an ODMTS service and optimal paths for each user under the current design by considering both their duration and cost; the potential riders then evaluate the convenience of these paths, and decide whether to adopt the ODMTS service. It is important to note that this is a fundamental distinction between ODMTS and conventional transit systems, as the transit agencies provide the designated paths for riders by using on-demand shuttles and transit network, and potential riders make their adoption decisions accordingly.
The remainder of this section is structured as follows: Sections 5.15.3 study ODMTS-DA problem using the bilevel optimization framework and explain the employment of heuristic algorithms to solve ODMTS-DA. Section 5.4 provides an overview of the three real-world case studies, which are categorized by size as medium, large, and extra-large. Finally, Sections 5.55.7 present the results over the corresponding three cases.
5.1 ODMTS-DA Bilevel Formulation
The analogy between ODMTS-DA and the generalized framework TN-DA presented in Section 3 can be summarized as follows:
  • Investment Function Invest: Sums of cost of opening each design arc, where the cost of opening $z_{hl}$ is denoted as $\beta_{hl}$. The unit of $\beta_{hl}$ should be consistent with $g^{r}$ and $\varphi^{r}$.
  • Subproblem $\text{\sc OPT-PATH}^{r}$: Finds the shortest path $\pi^{r}$ for trip $r$ from its origin to destination in terms of the cost $g^{r}$. It is important to emphasize that this subproblem aims to emulate the functionality of a rider mobile app (a realistic rider app example can be found in the Appendix of Van Hentenryck et al. (2023)), suggesting an ODMTS path by considering a convex combination of the agency’s monetary cost and the rider’s travel time. As a result, the solution to $\text{\sc OPT-PATH}^{r}$ is not necessarily the shortest path solely in terms of travel duration for trip $r$. Furthermore, the path can be multimodal, such as shuttle $\rightarrow$ bus $\rightarrow$ shuttle, or simply a single shuttle leg, namely direct shuttle. However, ODMTS does not allow a path with only two shuttle legs involved, i.e., shuttle $\rightarrow$ shuttle. Note that the existence of direct shuttle trips will guarantee the subproblem is always feasible.
  • Function PATH-COST: Computes the weighted cost $g^{r}$ of path $\pi^{r}$, based on the travel time and on-demand shuttle operating cost.
  • Function PATH-REVENUE: Returns a constant revenue value $\varphi^{r}$, because the ticket price is the same for all paths in ODMTS in practice.
The detailed bilevel formulation of ODMTS-DA can be found in Appendix 9.
5.2 Existing Methods for Solving ODMTS-DA
To solve the ODMTS-DA problem, Basciftci and Van Hentenryck (2023) proposed an exact decomposition algorithm whose resulting design can be considered as the equilibrium point of the game between the transit agency and the potential transit users. The effectiveness of this exact algorithm was demonstrated in a case study conducted in Ann Arbor and Ypsilanti area in Michigan, USA. Nevertheless, this exact algorithm becomes computationally challenging on large-scale instances. Furthermore, due to the complexity of the exact algorithm, it is even practically challenging to obtain high-quality sub-optimal solutions on large-scale problems, demonstrating the need to develop effective heuristics for providing fast approximations to the optimal solutions of large-scale instances with desired adoption properties of the resulting network designs.
As an alternative approach, another method, known as P-Path, reformulates the corresponding bilevel problem into a single-level Mixed-Integer Programming (MIP) problem (Guan et al. 2024). P-Path is capable of finding optimal solutions efficiently for medium-sized ODMTS-DA instances and within a reasonable time frame for large-scale instances. However, its computational resource requirements (e.g., memory) can be substantial because of the size of the resulting MIPs. Consequently, heuristic algorithms are beneficial, as they not only provide computational efficiency but might also offer a high-quality initial solution for P-Path.
5.3 Heuristic Algorithms for ODMTS-DA
This section illustrates how heuristic algorithms can be utilized to solve the ODMTS-DA problem and presents additional technical details to be utilized as part of the arc-based algorithms tailored to this problem setting. Building upon the general concepts discussed in Section 4, the heuristic algorithms make use of the ODMTS Design with Fixed Demand (ODMTS-DFD) problem and the choice model separately. The ODMTS-DFD problem is a specific form of the TN-DFD problem that is tailored to the ODMTS settings. The full formulation for ODMTS-DFD can be found in Appendix 9. To evaluate a design found by ODMTS-DFD, the evaluation function $\text{\sc Eval}(\mathbf{z})$ is adopted from the objective function of ODMTS-DA where $\pi^{r}$ and $\delta^{r}$ can all be deduced from $\mathbf{z}$, and the revenue term $\varphi^{r}$ for core trips are dropped because it remains as a constant value:
$\text{\sc Eval-ODMTS}(\mathbf{z})=\sum_{h,l\in H}\beta_{hl}z_{hl}+\sum_{r\in T_{core}}p^{r}g_{\pi}^{r}+\sum_{r\in T_{latent}}p^{r}\delta^{r}(g_{\pi}^{r}-\varphi^{r})$ (7)
[…]
For arc-based algorithms—Algorithms arc-S1 and arc-S2, the set of all candidate sub-networks $D^{k}$ at each iteration $k$ is restricted to include only directed bus cycles. Given that $\textbf{z}_{temp}^{k}$, $\textbf{z}_{unfixed}^{k}$, and $\textbf{z}_{fixed}$ all satisfy constraint (10b), the algorithm of Johnson (1975) can be used to find the set $D^{k}$ of directed cycles (lines 6 and 7 in Algorithms arc-S1 and arc-S2, respectively). The arc-based algorithms have been tested with a number of expansion rules to construct the Expand based on the specific nature of ODMTS. The expansion rules used at iteration $k$ includes:
  1. (a)
    All latent trips that adopt $\textbf{z}^{k}$. This is the Full Adoption Expansion Rule.
  2. (b)
    All latent trips that adopt $\textbf{z}^{k}$ and are not served by direct shuttle, i.e., they use fixed-routes.
  3. (c)
    All latent trips that adopt $\textbf{z}^{k}$ and satisfies $\text{UB}^{r}\leq\alpha^{r}\cdot t_{cur}^{r}$, where the term $\text{UB}^{r}$ is an upper bound on ODMTS travel time. The derivation of this bound can be found in Appendix 9.
The definition of rule (c) may seem overly complex; however, it leads to a useful proposition presented below, with its proof also provided in Appendix 9.
Proposition 5
When using Time-Based as Choice Model $\mathcal{C}^{r}$ (Equation (3)) in ODMTS-DA, the ODMTS design found by Algorithm arc-S1 with expansion rule (c) satisfies the Correct Adoption property (Property 2), i.e., $\%A_{false}=0$.
Additionally, when applying Algorithm arc-S2 in practice, it is worth noticing that the first rule should not be rule (a) as this rule includes all adopting latent trips, covering the trips considered in other expansion rules.
5.4 Overview on ODMTS-DA Case Studies
This section summarizes the three ODMTS-DA case studies. Table 2 provides a comparison of the settings across these cases by providing instance and solution approach information.
Size Location $|N|$ $|H|$ $|Z|$ $|T_{core}|$ $|T_{latent}|$ Congestion Choice Model $\mathcal{C}^{r}$ Methods
Medium Ypsilanti 1267 10 90 937 566 No Time-Based Exact Alg., P-Path, Heuristics
Large Atlanta 2426 58 828 15,478 36,283 No Time-Transfer-Based P-Path, Heuristics
Extra Large Atlanta 2426 66 2034 15,478 36,283 Yes Time-Based Exact Alg., Heuristics
Table 2: A summary on the three ODMTS-DA case studies
The first case study utilizes a real dataset from AAATA (https://www.theride.org Last Visited Date: September 30, 2025), the transit agency serving the greater Ypsilanti and Ann Arbor region of Michigan, USA. This case study assumes that the agency will replace the existing transit system entirely by implementing an ODMTS. Since the exact algorithm and P-Path can both successfully find the optimal design in a reasonable amount of time on this regular-sized dataset, this case study aims at demonstrating the workability of the heuristics proposed in Section 4 by comparing computational results obtained from different methods in terms of solution quality, adoption properties, and run time. Additional information on the dataset and experimental settings can be found in Appendix 11.
The second case study conducted with real travel data represents a regular workday in Atlanta, Georgia, to evaluate the performance of the heuristic algorithms on large-scale instances. This case study assumes that the agency will retain the rail system, remove the existing bus services, and integrate the rail system with newly designed ODMTS. As a result, all existing four rail lines are preserved, and all 38 rail stations are designated as ODMTS hubs (set $H$). Visualizations and data processing details are available in Appendix 12. As shown by Guan et al. (2024), P-Path can efficiently find the optimal solution for large-scale ODMTS-DA instances within a reasonable time frame and clearly outperforms the exact algorithm in computational efficiency. Therefore, in this case study, only P-Path and the heuristic algorithms are compared. Furthermore, to showcase that ODMTS-DA can take the choice model $\mathcal{C}^{r}$ as a black-box, another choice model, namely Time-Transfer-Based, is applied. In this choice model, a path $\pi$ is evaluated based on both travel duration and the number of required transfers, which is denoted by $l_{\pi}^{r}$. For instance, a “shuttle $\rightarrow$ rail $\rightarrow$ bus” path involves 2 transfers. Riders also have a transfer tolerance ($l_{ub}^{r}$), which is typically a low number and is set to 3 for all riders in this case study. The $\alpha$ value for all latent riders is set to 1.5.
$\text{\sc Time-Transfer-Based}:\mathcal{C}^{r}(\pi)\equiv\mathbbm{1}(t_{\pi}^{r}\leq\alpha^{r}\ t^{r}_{cur})\land\mathbbm{1}(l_{\pi}^{r}\leq l_{ub}^{r})$ (8)
The third case study uses the same data as the second case study, however it incorporates a larger set for $Z$ by considering more arcs between hubs (see details in Appendix 12). Preliminary investigation indicates that P-Path cannot handle this setting due to its massive memory requirement, exceeding 500 GB with commercial solver Gurobi 9.5. However, the exact algorithm, which is based on a decomposition approach, can partially solve this instance. Therefore, in this case, the exact algorithm is chosen as the baseline for evaluating the heuristic algorithms. The choice model is switched back to Time-Based to showcase the connection with Proposition 5. Additionally, this case study incorporates congestion modeling to enhance its realism.
The design objectives reported in the following subsections are based on Eval-ODMTS. Note that, although only a subset of trips is used to obtain the design in all algorithms, the design evaluation is computed with all core trips and latent trips to account for all riders considered.
5.5 Results: Medium Instance in Ypsilanti
Table 3 compares the computational performances of all solution methods along with the quality of their resulting network designs. The column ‘# Total Itr.’ reports the number of iterations, corresponding to the number of ODMTS-DFD problems solved during the runs. The column ‘# Outer Itr.’ indicates the number of iterations for $\rho$-GAGR, where its corresponding total iterations represent the number of iterations of $\eta$-GRRE (its subproblem). First, it can be seen that Algorithm $\rho$-GAGR requires significantly more computational time, whereas Algorithm $\eta$-GRRE takes about the same amount of computational time as Algorithm $\rho$-GRAD. Note that a practical approach to reduce the running time is to increase the step sizes of the trip-based algorithms. On the other hand, arc-based algorithms take less time compared to the trip-based algorithms. Secondly, four out of seven heuristic runs find the optimal solution, whereas Algorithm arc-S1 (with both rule (a) and rule (c)) and Algorithm arc-S2 (with rules (c) and (a)) report solutions with small optimality gaps. Despite having the same gap, the two arc-based algorithms report different results, due to their different expansion rules. Moreover, Propositions 2, 4, and 5 can be related to the results on False Rejection Rate and False Adoption Rate. In general, for transit agencies, small $\%R_{false}$ or $\%A_{false}$ values are indications of advantageous ODMTS designs when the optimal solution is unavailable. Three additional experiments with different parameter settings are provided in Appendix 11, demonstrating the performances of the proposed algorithms while providing network designs with desired adoption properties.
Method Step Size Trip Rule # Outer Itr. # Total Itr. Run Time (min) Design Obj % Opt Gap. % Gap $\%R_{false}$ $\%A_{false}$
Exact Alg. - - - - 360.29 29962.79 1.85 - 0.00 0.00
P-Path - - - - 3.34 29962.79 - 0.00 0.00 0.00
Alg. $\rho$-GRAD 10 - 1 26 2.48 29962.79 - 0.00 0.00 0.53
Alg. $\eta$-GRRE 10 - 1 26 2.49 29962.79 - 0.00 11.13 0.53
Alg. $\rho$-GAGR 10 - 26 357 47.14 29962.79 - 0.00 0.00 1.41
Alg. arc-S1 - (a) 1 8 1.05 30019.14 - 0.19 0.00 34.28
Alg. arc-S1 - (c) 1 7 1.13 30020.26 - 0.19 13.78 0.00
Alg. arc-S2 - (b)(a) 1 8 1.06 29962.79 - 0.00 0.00 2.65
Alg. arc-S2 - (c)(a) 1 8 0.95 30020.26 - 0.19 0.00 0.00
Table 3: Performance comparison on the medium instance. The exact algorithm reports optimality gap after 6 hours. The heuristics report the gap between themselves and the optimal solution (P-Path).
5.6 Results: Large Instance in Atlanta
Table 4 presents the results from heuristic algorithms on a large-scale instance, using the optimal solution from the P-Path model as a benchmark. For trip-based iterative algorithms, three step sizes, i.e., 1000, 2000, and 3000, were tested, and the step size associated with the best design objective is reported. The results show that while the heuristics do not reach the optimal solution, they still produce high-quality designs in a very short amount of time, whereas the optimal solution requires nearly five days of computation. Note that due to the structure of the ODMTS-DFD problem, the heuristics can utilize multi-threading techniques; here, 12 threads are employed. Additionally, using the solution from Algorithm arc-S2 as a warm-start for P-Path (because it provides the best solution in a shorter amount of time) can help reduce overall computation time. However, as pointed out by Guan et al. (2024), starting P-Path with a heuristic solution does not necessarily guarantee a faster path to the optimal solution.
One may wonder why heuristics are necessary if an optimal solution is available for instances up to a certain size, even though it is time-consuming to obtain. One of the reasons is that a transit agency may need to quickly develop a transit plan for special events, and not all locations have access to powerful computing resources capable of solving a large-scale P-Path model. In these circumstances, the heuristics can be useful. Furthermore, as noted by Guan et al. (2024), P-Path may spend a substantial amount of computing time on closing the final 1% of the optimality gap during the branch-and-bound processes. The solution reported here required 73% of its total computation time (around 85 out of the 116.48 hours) to address this last 1%, while the heuristics can achieve a better solution within just a few minutes.
Method Best Step Size Trip Rule # Outer Itr. # Total Itr. Run. Time (hour) Design Obj %Gap $\%R_{false}$ $\%A_{false}$
P-Path - - - - 116.48 214463.71 0.00 0.00 0.00
P-Path Warm - - - - 88.19 214463.71 0.00 0.00 0.00
Alg. $\rho$-GRAD 3000 - 1 8 0.07 214906.96 0.21 0.00 7.90
Alg. $\eta$-GRRE 1000 - 1 17 0.15 221265.96 3.17 4.72 0.84
Alg. $\rho$-GAGR 2000 - 10 69 0.66 214906.96 0.21 0.40 6.70
Alg. arc-S1 - (a) 1 20 0.72 214846.58 0.18 0.00 14.37
Alg. arc-S1 - (c) 1 13 0.31 242229.26 12.95 31.43 0.00
Alg. arc-S2 - (b),(a) 1 22 0.60 214846.58 0.18 0.00 7.29
Alg. arc-S2 - (c),(a) 1 22 0.50 214906.96 0.21 0.00 11.60
Table 4: A comparison between P-Path and the heuristics on the large instance. P-Path is solved to optimal. Warm stands for warm-start, the reported time includes both the heuristic and the P-Path computation time.
5.7 Results: Extra-Large Instance in Atlanta
This section evaluates the computational results of the extra-large case study from various viewpoints along with the analyses on the resulting network designs. Additional discussions on results such as adoptions, operating cost, travel time, and travel distance can be found in Appendix 13.
Algorithmic Performances
Table 5 compares the performances of all algorithms. The most obvious finding to emerge from the analysis is that the heuristic algorithms can outperform the exact algorithm in a reasonable amount of time. Moreover, after 48 hours, the exact algorithm finds a solution with objective 213828.87, which is only 1.61% better than the 24-hour running time result. Thus, the proposed iterative algorithms can efficiently provide high-quality solutions to the ODMTS-DA problem. The only exception is Algorithm arc-S1 with rule (c), whose greater objective value can be attributed to the strict nature of rule (c).
Method Best Step Size Trip Rule # Outer Itr. # Total Itr. Run. Time (hour) Design Obj. % Gap $\%R_{false}$ $\%A_{false}$
Exact Alg. - - - - 26.20 220557.59 - - -
Alg. $\rho$-GRAD 3000 - 1 9 1.38 209548.18 -4.99 0.00 7.18
Alg. $\eta$-GRRE 3000 - 1 11 1.66 213409.00 -3.24 5.46 0.49
Alg. $\rho$-GAGR 3000 - 8 53 9.72 209412.04 -5.05 1.41 4.07
Alg. arc-S1 - (a) 1 20 5.58 208664.07 -5.39 0.00 20.16
Alg. arc-S1 - (c) 1 11 0.74 246561.82 11.79 63.78 0.00
Alg. arc-S2 - (b), (a) 1 25 3.69 211253.00 -4.22 0.00 11.48
Alg. arc-S2 - (c), (a) 1 27 4.75 208294.23 -5.56 0.00 18.48
Table 5: Performance comparison between all algorithms on the extra-large instance. The exact algorithm reports the best feasible solution within a 24 hours running-time limit. ’% Gap’ shows the gap between the objective value reported by the heuristics and the objective reported by the exact algorithm.
The table also shows that Algorithms $\rho$-GRAD and $\eta$-GRRE achieve the second and third fastest termination among all iterative algorithms, although they fail to discover a better solution compared to the other approaches. These performances are likely to be related to the adoption set $C$ of Algorithm $\rho$-GRAD and the rejecting set $R$ of Algorithm $\eta$-GRRE. These sets are designed to speed up the design searching processes, yet it leads the algorithm to overlook multiple intermediate ODMTS designs. Note that multi-threading is disabled for this instance in order to produce a fair comparison on all algorithms. However, when the multi-threading techniques for subproblems in ODMTS-DFD are allowed, the heuristics can be massively sped up. For instance, with 18 threads and a step size of 3000, Algorithm $\eta$-GRRE can terminate in 0.24 hours. On the other hand, the combined trip-based iterative algorithm can indeed explore more ODMTS designs and provide a lower objective: algorithm $\rho$-GAGR reports a better objective than Algorithm $\eta$-GRRE. However, it does not outperform Algorithm $\rho$-GRAD. A possible explanation for this result can be the choices of step sizes.
In this particular case study, Algorithm $\rho$-GRAD produces a better balance between running time and design objective among all the trip-based iterative algorithms. The reported $\%R_{false}$ value is consistent with Proposition 2. In other words, with ODMTS design with $\%R_{false}=0$, all riders excluded during the designing procedure reject the system. However, there will be $\%A_{false}$ riders who reject the system even though they were included in the design procedure, e.g., 7.18% latent trips reject the design when using Algorithm $\rho$-GRAD although they are in $\hat{T}$. Thus, this type of ODMTS designs can be treated as an upper limit when agencies plan to expand their transit network to serve more passengers.
For the arc-based iterative algorithms, Algorithm arc-S1 has consistent $\%R_{false}$ and $\%A_{false}$ values with Propositions 4 and 5. Algorithm arc-S1 with rule (c) is the only case that guarantees $\%A_{false}=0$ in this study. That is, all riders used to design the ODMTS service will adopt it. Although this is a relatively conservative approach, the transit agency can start to expand the services based on this ODMTS design. Secondly, the extended version of Algorithm arc-S1, i.e., Algorithm arc-S2, also found high-quality solutions with a considerable drop in running time and more iterations compared to Algorithm arc-S1 with trip rule (a). This result can be explained by the fact that the expansion rule of Algorithm arc-S1 is significantly looser than the first rule employed by Algorithm arc-S2. Thus, the size of the $\hat{T}$ in Algorithm arc-S2 for each iteration is substantially smaller, resulting in shorter running time per iteration. Closer inspection of the table shows that the solution found by Algorithm arc-S2 (with rules (c) and (a)) outperforms all other algorithms and successfully establishes the smallest upper bound on the design objective in a relatively short duration. Furthermore, the results found by Algorithm arc-S2 are also consistent with Corollary 2.
In summary, the results compare the algorithmic performances of all heuristic algorithms. A few practical recommendations for choosing the algorithms might then be summarized: (i) use Algorithms $\rho$-GRAD or $\eta$-GRRE with relatively large step size to conduct a fast exploration process, (ii) employ Algorithm arc-S1 or Algorithm arc-S2 with certain trip expansion rules in order to build the upper bound of the problem, and (iii) apply either Algorithm $\rho$-GRAD or Algorithm arc-S2 when the balance between solution quality and running time is demanded.
ODMTS Designs
(a) Only with Core Trips
(b) Exact Alg. in 24 hours
(c) Alg. $\rho$-GRAD
(d) Alg. $\eta$-GRRE
(e) Alg. $\rho$-GAGR
(f) Alg. arc-S1 rule (a)
(g) Alg. arc-S1 rule (c)
(h) Alg. arc-S2 rules (b) (a)
(i) Alg. arc-S2 rules (c) (a)
Figure 4: The ODMTS designs found by the Algorithms for the extra-large case study. Black polylines and green straight lines are utilized to represent opened bus arcs and opened non-direct shuttle legs, respectively. Thickness of the black polylines indicates the number of users using the arcs. The rail systems are in red, yellow, dark green, and blue.
The ODMTS designs reported by the exact algorithm and all iterative algorithms are illustrated in Figure 4. In particular, Figure 4(a) presents the design when solely using the core trip set $T_{core}$, i.e., $\hat{T}=T_{core}$. It is thus the ODMTS design discovered by all trip-based heuristic algorithms at the first iteration. The other figures show the ODMTS designs obtained from the algorithms. The non-direct shuttle legs are on-demand shuttle services that connect a hub and a rider’s origin stop or destination stop. For clearer interpretation of the maps, the main concentrations of latent trips in this case study are in the East Atlanta region (east of the red and yellow rail lines and north of the blue rail line), which is also where the primary differences in the designs appear. Additional details can be found in the latent demand heatmaps provided in Appendix 12.2. Compared to the design in Figure 4(a), it can be seen that additional trips in $\hat{T}$ generally lead to a more complex network in the East Atlanta region. Closer inspection of all ODMTS designs show that the majority of bus arcs are used to deliver riders to the rail system. This might be explained by the fact that the rail system operates at a high frequency and connects the major districts such as Downtown (where all 4 rail lines intersect), Midtown (just north of Downtown, where the red and yellow lines overlap), and the city of Sandy Springs (the northernmost section of the red line). The simplest design among Figure 4 is found by Algorithm arc-S1 with rule (c), illustrated in Figure 4(g), which is identical to the design with only core trips as input. The designs found by arc-based algorithms with looser expansion rule can be seen in Figures 4(f)4(h), and 4(i), and these designs are relatively more convoluted as expected. Additionally, Figure 4(b) presents the design reported by the exact algorithm after 24 hours running time. Comparing to the other designs, this design shows a few differences in the areas with latent riders, e.g., in the South and Northwest regions. This can be explained by the 1% or 8 hours terminating criteria for the master problem in the ODMTS-DA problem (Figure 6).
6 TN-DA Example 2: SCTS Design with Adoptions Problem
In order to show that TN-DA can be tailored for different transit network design problems, this section introduces another example, namely Scooters-Connected Transit Systems Design with Adoptions (SCTS-DA). As pointed out by McQueen and Clifton (2022) and Yan et al. (2023), for connecting to transit lines, e-scooters started to emerge as a solution to the first and last mile challenges. Different to the ODMTS example, the SCTS-DA problem assumes there are two independent service operators involved: (1) a transit agency responsible for operating high-frequency buses and rails, and (2) a Transportation Network Company (TNC) that offers e-scooter services. This problem assumes a relatively simplified scenario where there is a sufficient number of e-scooters available throughout the city, and the TNC is responsible for managing and relocating the e-scooters as needed. In this context, the primary goal of the problem is to design an efficient high-frequency bus network that can be seamlessly connected with the e-scooter services while taking into account rider adoption considerations.
6.1 SCTS-DA Bilevel Formulation
The analogy between SCTS-DA and the generalized framework TN-DA, can be summarized as follows:
  • Investment Function Invest: Sums of cost of opening each design arc, where the cost of opening $z_{hl}$ is denoted as $\beta_{hl}$. The unit of $\beta_{hl}$ should be consistent with $g^{r}$ and $\varphi^{r}$.
  • Subproblem $\text{\sc OPT-PATH}^{r}$: Finds the shortest path $\pi^{r}$ for trip $r$ in terms of the cost $g^{r}$. Assume that the riders’ primary goal is minimizing travel time; this is equivalent to the conventional shortest path based on travel time. A direct scooter trip between the origin and destination ensures that the subproblem is feasible.
  • Function PATH-COST: Computes the weighted cost $g^{r}$ of path $\pi^{r}$. The cost $g^{r}$ is weighted based on $\pi^{r}$’s travel time of path $t_{\pi}^{r}$.
  • Function PATH-REVENUE: Returns constant revenue value $\varphi^{r}$ because transit ticket price is flat for all users. Scooter fees are not considered as part of the revenue as the scooter services are not offered by the transit agency.
The detailed bilevel formulation of SCTS-DA can be found in Appendix 10.
The decision model employed in this case study is referred to as Time-Based-Scooter (Equation (9)). It takes into account three factors: firstly, the travel duration within the transit system is required to be smaller than or equal to an $\alpha^{r}$ value multiplied by the travel duration of the current mode of transportation. Secondly, riders will only adopt the transit system if they actually utilize buses or trains, which will incur income for the transit agency. Lastly, riders only use the transit system if they travel on scooter less than a threshold value $d_{ub}^{r}$ given that scooters are not an efficient mode for longer distances. In the following equation, $x_{hl}$ and $y_{ij}$ denote if trip $r$ use design arc $(h,l)$ and scooter arc $(i,j)$. The travel distance between locations $i$ and $j$ is $d_{ij}$.
$\text{{\sc Time-Based-Scooter}}:{\cal C}^{r}(\pi)\equiv\mathbbm{1}(t_{\pi}^{r}\leq\alpha^{r}\ t^{r}_{cur})\land\mathbbm{1}\bigg(\sum\limits_{h,l\in H}x_{hl}^{r}>0\bigg)\land\mathbbm{1}\bigg(\sum\limits_{i,j\in N}y_{ij}^{r}\cdot d_{ij}\leq d_{ub}^{r}\bigg)$ (9)
6.2 Heuristic Algorithms for SCTS-DA
This section presents the heuristic settings for SCTS-DA; the formulations and the evaluation function are documented in Appendix 10.
In the case of the trip-based iterative algorithms, the value used to rank trips, denoted as $v^{r}$, is determined based on the time traveled by e-scooters since e-scooter services typically charge riders based on the amount of time they utilize the scooters. For arc-based algorithms, the expansion rules used at iteration $k$ include:
  1. (i)
    All latent trips that adopt $\textbf{z}^{k}$. This is the Full Adoption Expansion Rule.
  2. (ii)
    All latent trips that adopt $\textbf{z}^{k}$ and use the backbone transit lines. Backbone refers to arcs that are preserved in the system without requiring design through solving TN-DA, e.g., rail system in Atlanta.
The first rule considers an expansive case by including all latent riders adopting, whereas second rule considers a subset of these trips by reflecting a common scenario where scooters are used for first-and last-mile connections to primary transit corridors, typically rail lines. For instance, these transit stations are usually equipped with convenient parking areas for scooters in Atlanta.
6.3 Results: SCTS-DA in Atlanta
Similar to the large-scale ODMTS-DA case studies explained in the previous section for Atlanta, this case study employs the same dataset and parameters for buses and rails (see Appendix 12.5). For scooters, the average speed is assumed to be $16.1\text{\,}\mathrm{k}\mathrm{m}\mathrm{/}\mathrm{h}$ ($10\text{\,}\mathrm{m}\mathrm{p}\mathrm{h}$) when traveling in Atlanta, and the scooter distance bound $d_{ub}^{r}$ is set at $24\text{\,}\mathrm{k}\mathrm{m}$ ($15\text{\,}\mathrm{m}\mathrm{i}\mathrm{l}\mathrm{e}\mathrm{s}$) for all trips. Moreover, this case study also employs the congested matrices for cars and buses. Due to the similar structure of this problem to ODMTS-DA, a modified version P-Path is derived to seek the optimal solution of SCTS-DA, and treated as a benchmark for this setting.
The computational results are reported in Table 6. Here, the two arc-based algorithms can find a solution that have zero False Rejection Rate and False Adoption Rate, which are the desired properties of an optimal solution. Although Algorithms $\rho$-GRAD and $\eta$-GRRE find the solution with higher objective, they terminate in a short duration. On the other hand, Algorithm $\rho$-GAGR is ineffective in producing advantageous results for this case study. In Figure 5, it can be observed that the designs obtained by the three trip-based heuristic algorithms (Figure 5(a)) are identical, as do the designs discovered by the two arc-based algorithms (Figure 5(b)). Furthermore, it is worth mentioning that the designs resulting from the trip-based algorithms are quite similar to the design derived exclusively from core trips.
(a) Trip-Based Algorithms
(b) Arc-Based Algorithms and Modified P-Path
Figure 5: The Designs Found by the Heuristic Algorithms and modified P-Path for the SCTS-DA problem in Atlanta. Light green lines represent scooter connections.
Method step size trip rule # outer itr. # total itr. time (hours) design obj % Gap $\%R_{false}$ $\%A_{false}$
Modified P-Path - - - - 4.58 1051031.55 0.00 0.00 0.00
Alg. $\rho$-GRAD 500 - 1 12 0.11 1055265.30 0.40 11.61 0.00
Alg. $\eta$-GRRE 500 - 1 12 0.10 1055265.30 0.40 11.61 0.00
Alg. $\rho$-GAGR 500 - 12 79 0.70 1055265.30 0.40 11.61 0.00
Alg. arc-S1 - (i) 1 15 0.55 1051031.55 0.00 0.00 0.00
Alg. arc-S2 - (ii), (i) 1 16 0.57 1051031.55 0.00 0.00 0.00
Table 6: Performance comparison between all algorithms on the Large-scale experiment.
7 Conclusion
This study aims at capturing the latent demand for the transit agency such that the transit services can concurrently provide high-quality services to existing riders and extend the services to current underserved areas. The decision-making processes of both the transit agency and the latent riders form a game-like interaction. To model this dynamic interaction, this paper first introduces a bilevel optimization framework called Transit Network Design with Adoptions (TN-DA). The framework essentially models a network design problem with an incorporated black-box-like mode choice model. Achieving success in solving the TN-DA can identify the equilibrium point between the agency and the latent riders. However, even with an exact solution algorithm, due to the bilevel and combinatorial nature of the TN-DA, practically solving it can lead to computational challenges, particularly for large-scale instances.
This paper then attempts to address this limitation and proposes five heuristic algorithms to efficiently approximate the optimal solution of the TN-DA problem. The fundamental idea of these iterative algorithms is to separate the previous bilevel framework into two individual but simpler components, i.e., a regular network design problem with no adoption awareness (the TN-DFD problem in Figure 2) and an evaluation procedure based on a mode choice function. These heuristic algorithms can be further categorized into two general classes based on their algorithmic designs: (1) trip-based algorithms and (2) arc-based algorithms. In general, at each iteration, the trip-based algorithms reconstruct the trip sets for the TN-DFD problem and the arc-based algorithms attempt to fix a certain amount of bus arcs. The advantages and limitations of each algorithm are discussed in this paper, and several theoretical properties are derived from the optimal solutions in order to provide more guidelines for the iterative algorithms under varied circumstances. In particular, two key properties, correct rejection (Property 1) and correct adoption (Property 2), are introduced. Then, the false rejection rate and the false adoption rate are proposed as metrics to evaluate the ODMTS designs, providing additional insights on the performance of the heuristic algorithms when an optimal solution becomes unavailable for large-scale instances.
This paper further presents comprehensive case studies on two different transit systems to validate the heuristics. For the first transit system—On-Demand Multimodal Transit Systems (ODMTS), one case study is carried out with a normal-sized transit dataset collected by the transit agency in the Ann Arbor and Ypsilanti area in the state of Michigan. The results demonstrate that the heuristic algorithms can discover the optimal ODMTS design. Another case study is based on multi-sourced data and includes a very large numbers of stops and trips in Atlanta in the state of Georgia. The main goal of this case study is to show the advantages of the heuristic algorithms on large-scale instances. Compared to the exact algorithm, the heuristic algorithms achieve better objective values in considerably less amount of time. The second transit system—Scooters-Connected Transit Systems (SCTS) are discussed in Section 6. These case studies serve as evidence that the TN-DA framework can be readily tailored to various transit systems and exhibit strong performance on large-scale scenarios while providing practical insights.
[…]
This work was partially supported by NSF Leap-HI Grant No. 1854684, NSF Grant No. 2112533, NSF CMMI Grant No. 2434302, and T-SCORE Grant 69A3552047141 from the Department of Transportation.